# The Pitfalls of Next-Token Prediction

### v2

@@ Abstract
Can a mere next-token predictor faithfully model human intelligence? We crystallize this emerging concern and correct popular misconceptions surrounding it, and advocate a simple multi-token objective. As a starting point, we argue that the two often-conflated phases of next-token prediction — autoregressive inference and teacher-forced training — must be treated distinctly. The popular criticism that errors can compound during autoregressive inference, crucially assumes that teacher-forcing has learned an accurate next-token predictor. This assumption sidesteps a more deep-rooted problem we expose: in certain classes of tasks, teacher-forcing can simply fail to learn an accurate next-token predictor in the first place. We describe a general mechanism of how teacher-forcing can fail, and design a minimal planning task where both the Transformer and the Mamba architecture empirically fail in that manner — remarkably, despite the task being straightforward to learn. Finally, we provide preliminary evidence that this failure can be resolved using a simple modification that predicts multiple tokens in advance. We hope this finding can ground future debates and inspire explorations beyond the next-token prediction paradigm. We make our code available under https://github.com/gregorbachmann/Next-Token-Failures
@@ 1 Introduction
Although this argument is compelling, a pessimist would worry, realistically, even with minor imperfections in the next-token predictor, the accuracy may fall spectacularly for long sequences (Kääriäinen, 2006; Ross & Bagnell, 2010; LeCun, 2024; Dziri et al., 2024). Say, even if every next-token error is as little as $0.01$, the probability of encountering an erroneous token exponentially compounds along the way, and by the end of $200$ tokens, blows up to $0.86$.
@@
Empirically, we demonstrate that the above mechanism leads to complete in-distribution failure in a path-finding setup on a graph, that we propose as a minimal lookahead task. We design our setup in a way that it is demonstrably straightforward to solve that the failure of any model is remarkable. Yet, we observe failure for both the Transformer (Vaswani et al., 2017) and the Mamba architecture, a structured state space model (Gu & Dao, 2023). We also find that a form of teacherless training that predicts multiple future tokens (Monea et al., 2023) is (in some settings) able to circumvent this failure. Thus, we pinpoint a precise and easy-to-learn scenario where, rather than properties that are criticized in existing literature — like convolution or recurrence or autoregressive inference (see §6), — it is next-token prediction during training that is at fault.
@@ 4.1 Path-Finding on Path-Star Graphs: A Minimal and Easy Lookahead Task
Consider a path-finding problem on a directed graph $G$ with a set of nodes $\{{v}_{\tt{start}},{v}_{\tt{goal}},v_{1},v_{2},\ldots\}$. The graph is a “path-star” graph with ${v}_{\tt{start}}$ as the central node, with at least $2$ paths (each of length $l\geq 2$ edges) emanating from it, with a unique path ending in ${v}_{\tt{goal}}$. The task is to find a path from ${v}_{\tt{start}}$ to ${v}_{\tt{goal}}$. Correspondingly, we assume that the distribution $\mathcal{D}$ is over sequences where the prefix $\bm{p}$ represents a (randomly generated) graph, and the response represents the path from the start to the goal. In particular, we sample a graph $G$ which is represented as an adjacency list as $\texttt{adj}(G)=e_{1},e_{2},\ldots$ where each edge $e=(v,v^{\prime})$ is represented such that $v^{\prime}$ farther away from ${v}_{\tt{start}}$ than $v$. We then set the prefix as $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$ so the model knows what the graph, and the desired start and goal states are. The ground truth response $\bm{r}$ corresponds to the sequence of vertices $\bm{r}={v}_{\tt{start}},\ldots{v}_{\tt{goal}}$ on the start-to-goal path. We visualize this construction in Fig. 1.
@@ 4.3 The Clever Hans Cheat
To illustrate this in our path-star example, without loss of generality, consider a ground truth path that is of the form $\bm{r}={v}_{\tt{start}},v_{1},v_{2},\ldots,{v}_{\tt{goal}}$. With a slight abuse of the indexing notation, let $\bm{r}_{<i}={v}_{\tt{start}},v_{1},\ldots,v_{i-1}$ be the prefix of length $i$ (so we index from 00 instead of $1$). Observe that nodes from $v_{2}$ onwards, until before ${v}_{\tt{goal}}$, have precisely one edge going “away” from ${v}_{\tt{start}}$. Thus, consider when the model is given as input, $(\bm{p},\bm{r}_{<i})$ where $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$, to fit the target $v_{i}$. The model first merely needs to scan the adjacency list $\texttt{adj}(G)$ within $\bm{p}$ for the one edge containing $v_{i-1}$ in the first position. Then, the model only has to predict the other node on that edge as $v_{i}$. Note though, this cheat cannot work on fitting the target $v_{1}$ given the input $\bm{r}_{<1}={v}_{\tt{start}}$ since ${v}_{\tt{start}}$ has many outward edges — we will address this node in the next section. We illustrate this difference between $v_{1}$ and the remaining tokens as the “easy” vs. “difficult” tokens in Fig. 2.
Crucially, the above cheating mechanism for fitting the easy tokens does not require any lookahead. It is simple, and implementable by an induction head-like module (Olsson et al., 2022). Owing to this simplicity, we hypothesize that these tokens will be quickly fit and ignored during training. This destroys useful signal for the model to efficiently learn the underlying “right-to-left” solution: the solution that requires looking at all tokens in $\bm{r}$, and then learning that they are simply the unique path from ${v}_{\tt{goal}}$ spelled in reverse.
We emphasize two key aspects of this cheating behavior. First, these shortcuts are unlike well-known shortcuts (see Remark 5) that map from the original input prefixes $\bm{p}$ to the ground truth $\bm{r}$. What we identify is unique to the mapping from the teacher-forced prefix $(\bm{p},\bm{r}_{<i})$ to $r_{i}$. We name this Clever Hans cheating. Another notable point is that this does not come from a dearth of samples: even if we had infinite training data at our disposal, the model can still fit the easy tokens of all that data by Clever Hans cheating.
@@ 5 Experimental Verification
For each experiment, we generate the training and test graphs from the same distribution $\mathcal{D}$, all with the same topology of ${G}_{d,l}(N)$ with fixed $d,l$ and $N$. Thus, any failure we demonstrate is an in-distribution failure, and does not arise from the inability to generalize to different problem lengths (Anil et al., 2022). We note that while the graphs are all of the same topology, this is not a trivial memorization problem for the model, since the graphs are labeled differently, and the adjacency list randomized — the model has to learn a general algorithm. Throughout the experiments, we fix the number of samples to $200k$ and fix the number of node values to $N=100$ across topologies to enable diverse instantiations of the topology for training and testing.
@@
We evaluate models from two architectural families to highlight that the failures are not tied to a particular architecture but stem from the next-token prediction objective. For Transformers, we use from-scratch GPT-Mini, and pretrained GPT-2 large (Radford et al., 2019). For recurrent models, we use from-scratch Mamba (Gu & Dao, 2023). We optimize using AdamW (Loshchilov & Hutter, 2019) until perfect training accuracy. To rule out grokking behaviour (Power et al., 2022), we trained the cheaper models for as long as $500$ epochs. More details are in §G.2.
@@
For a given distribution, we evaluate all our teacher-forced models by autoregressively generating solutions, and comparing that solution with the true one for an exact-match. We denote this accuracy as $\texttt{Acc}_{\texttt{ag}}(\texttt{LM}_{\theta})$ (see Eq 5) and report it for path-star graphs of varying topologies in Fig. 3 and Table 2. As observed, all models (even when pre-trained) struggle to learn the task accurately. The accuracies are precisely limited to the value when uniformly guessing a path from ${v}_{\tt{start}}$ i.e., $\approx\frac{1}{d}$, thus establishing complete in-distribution failure. This is so even when trained to fit sample sizes up to $200k$ to $100\%$ accuracy, and despite the fact that the training and test graphs have identical topology. Next, we quantitatively demonstrate how this stark failure arises from our two hypothesized mechanisms (Failure 2a2b).
@@
Recall that the Clever Hans cheat only applies to all but the first node $v_{1}$ after ${v}_{\tt{start}}$ lying on the path. After the Clever Hans cheat fits the rest of the path during training, we hypothesized that node $v_{1}$ may become impossible to learn since the model is deprived of all information about the subsequent targets. To quantify this behavior, we evaluate how well the model is able to predict the difficult first node, $v_{1}$. We measure this on the held-out set and denote this as $\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})$ (see Eq 7). As shown in Fig. 4 the model achieves a low $\texttt{Acc}_{\texttt{1st}}(\texttt{LM}_{\theta})$, approximately $1/d$. Thus, the model indeed fails to identify that $v_{1}$ is the one on the path to ${v}_{\tt{goal}}$ and instead randomly emitting one of the $d$ neighbors of ${v}_{\tt{start}}$.
Removing the Clever Hans cheat via teacherless training (Monea et al., 2023).
We now consider a training setup where we prevent Clever Hans cheating (Failure 2a) and examine how learning differs. Concretely, consider modifying teacher-forcing by replacing the input $\bm{r}$ (which reveals the ground truth) with an uninformative input $\bm{r}^{\$}$, consisting of the same special (“lookahead”) token \$currency-dollar\\$\$ repeated $l$ times. For supervision in the loss, we still use the original target $\bm{r}$. Thus, the model cannot fit the targets by looking at the prefixes $\bm{r}_{<i}$ and by predicting the next token $v_{i}$ via cheating. Instead, the model only has access to the graph description in $\bm{p}$ to lookahead and fit all the targets $v_{i}$ for $i=1,\ldots,l$. Formally, we maximize:
@@
We denote a model trained this way by $\texttt{LM}_{\theta}^{\$}$ and perform inference simply by conditioning on \$currency-dollar\\$\$ tokens i.e., to extract $\hat{r}_{i}$, we feed the uninformative prefix $\bm{r}^{\$}_{<i}$ as input, rather than autoregressively feeding the output $\hat{\bm{r}}_{<i}$ as input. We denote the resulting accuracy by $\texttt{Acc}_{\texttt{\$}}(\texttt{LM}_{\theta}^{\$})$ (see Eq 9). This training and inference setup was proposed in Monea et al. (2023) for the orthogonal goal of improving the computational efficiency of inference. Our goal however is to evaluate whether forcing the model to lookahead can dodge the Clever Hans cheat, thereby allowing the correct mechanism to be learned.
@@ 6 Related Work
Going beyond next-token prediction. Various works have explored architectures and objectives that train the backbone to go beyond next-token prediction. This includes non-autoregressive models (Gu et al., 2018), energy-based models (Dawid & LeCun, 2023), diffusion models (Gong et al., 2023), and variants of Transformers learning to either predict future tokens in a different ordering (Li et al., 2021) or all at one go (Qi et al., 2020; Monea et al., 2023), or injecting “lookahead” data (Du et al., 2023a). Teacherless training — proposed as “parallel speculative sampling (PaSS)” in Monea et al. (2023) — provides an arguably simple such approach that involves a trivial modification to teacher-forcing. Note that while research in parallel decoding too is concerned with predicting multiple future tokens (Stern et al., 2018), the goal is purely inference-time efficiency — which is also the setting under which Monea et al. (2023) propose teacherless training.
@@ Acknowledgments
We would like to thank Colin Raffel, Tiago Pimentel, and Surbhi Goel for their extensive feedback on a draft of this preprint, especially for pointers to some key references.
@@ Appendix B Teacher-Forcing Failure and Snowballing Failure are Distinct
Consider the model learned via teacher-forcing on the graph problem. During inference, we saw that it suffers a debilitating error right in the first step (with accuracy of $1/d$ for degree $d$ of the start node). Thus, during inference the error that is experienced is not from an accumulation over length. In fact, if only the first node is set correctly during inference, a model with the perfect Clever Hans cheat, would achieve $100\%$ accuracy rate. Such a model does not experience snowballing errors.
@@ Appendix D Other Remarks
      Node-ID-agnostic memorization: If the number of training data is $\Omega(|E|!)$, the model can still implement a form of memorization, but this requires a cleverer strategy that is agnostic to the node ID’s. Concretely, for a fixed assignment of node ID’s, there are only $O(|E|!)$ ways to permute the adjacency list. Assume that the model sees all such permutations during training (albeit with varying instantiations of the node IDs). During inference, the model can first look up whether the test input corresponds to an existing permutation it had witnessed (possibly with different node ID’s). For example $1\to 2;2\to 3;4\to 1$ would correspond to $10\to 20;20\to 30;40\to 10$. Then the model merely needs to recall from its memory, the index where the target token is located in the input for that permutation. The model can then look at the same index in the test input and output the node ID located there. Thus, if the target token was $4$ for the first sequence above, the model can output $40$ for the other sequence.
@@ E.4 Inference in Teacherless Training
In teacherless training, we make use of an uninformative input $\bm{r}^{\$}$ that simply corresponds to a series of dummy tokens denotes by \$currency-dollar\\$\$. During inference, instead of autoregressing on the model’s own output, we use this uninformative input. We formalize this below:
@@ F.1 Snowball Failure
[…]
If $\texttt{Acc}_{\texttt{sb}}(\texttt{LM}_{\theta})$ is $\approx 1$, then Failure 1 is not prominent in our task. If $\texttt{Acc}_{\texttt{sb}}(\texttt{LM}_{\theta})\ll 1$, then clearly teacher-forcing is responsible for surpressing errors in generation, strongly hinting at the fact that Failure 1 is at play. We display the results in Fig. 5 (left). We observe that the accuracy $\texttt{Acc}_{\texttt{sb}}$ is barely affected even for graphs with very long paths $L=40$.

### v3

@@ Abstract
Can a mere next-token predictor faithfully model human intelligence? We crystallize this emerging concern and correct popular misconceptions surrounding it, and advocate a simple multi-token objective. As a starting point, we argue that the two often-conflated phases of next-token prediction — autoregressive inference and teacher-forced training — must be treated distinctly. The popular criticism that errors can compound during autoregressive inference, crucially assumes that teacher-forcing has learned an accurate next-token predictor. This assumption sidesteps a more deep-rooted problem we expose: in certain classes of tasks, teacher-forcing can simply fail to learn an accurate next-token predictor in the first place. We describe a general mechanism of how teacher-forcing can fail, and design a minimal planning task where both the Transformer and the Mamba architecture empirically fail in that manner — remarkably, despite the task being straightforward to learn. Finally, we provide preliminary evidence that this failure can be resolved using teacherless training, a simple modification using dummy tokens that predicts multiple tokens in advance. We hope this finding can ground future debates and inspire explorations beyond the next-token prediction paradigm. We make our code available under https://github.com/gregorbachmann/Next-Token-Failures
@@ 1 Introduction
Although this argument is compelling, a pessimist would worry, realistically, even with minor imperfections in the next-token predictor, the accuracy may fall spectacularly for long sequences (Kääriäinen, 2006; Ross & Bagnell, 2010; LeCun, 2024; Dziri et al., 2024). Say, even if every next-token error is as little as 0.010.010.01, the probability of encountering an erroneous token exponentially compounds along the way, and by the end of 200200200 tokens, blows up to 0.860.860.86.
@@
Empirically, we demonstrate that the above mechanism leads to complete in-distribution failure in a path-finding setup on a graph, that we propose as a minimal lookahead task. We design our setup in a way that it is demonstrably straightforward to solve, implying that the failure of any model is remarkable. Yet, we observe failure for both the Transformer (Vaswani et al., 2017) and the Mamba architecture, a structured state space model (Gu & Dao, 2023). We then point towards a very simple multi-token modification called teacherless training — an idea that has appeared in other contexts (Tschannen et al., 2023; Monea et al., 2023; Zhao et al., 2023) — which predicts multiple future tokens and is able to circumvent this failure in some settings. Thus, we pinpoint a precise and easy-to-learn scenario where, rather than properties that are criticized in existing literature — like convolution or recurrence or autoregressive inference (see §6), — it is next-token prediction during training that is at fault.
@@ 4.1 Path-Finding on Path-Star Graphs: A Minimal and Easy Lookahead Task
Consider a path-finding problem on a directed graph $G$ with a set of nodes $\{{v}_{\tt{start}},{v}_{\tt{goal}},v_{1},v_{2},\ldots\}$. The graph is a “path-star” graph with ${v}_{\tt{start}}$ as the central node, with at least 222 paths (each of length $l\geq 2$ edges) emanating from it, with a unique path ending in ${v}_{\tt{goal}}$. The task is to find a path from ${v}_{\tt{start}}$ to ${v}_{\tt{goal}}$. Correspondingly, we assume that the distribution $\mathcal{D}$ is over sequences where the prefix $\bm{p}$ represents a (randomly generated) graph, and the response represents the path from the start to the goal. In particular, we sample a graph $G$ which is represented as an adjacency list as $\texttt{adj}(G)=e_{1},e_{2},\ldots$ where each edge $e=(v,v^{\prime})$ is represented such that $v^{\prime}$ farther away from ${v}_{\tt{start}}$ than $v$. We then set the prefix as $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$ so the model knows what the graph, and the desired start and goal states are. The ground truth response $\bm{r}$ corresponds to the sequence of vertices $\bm{r}={v}_{\tt{start}},\ldots{v}_{\tt{goal}}$ on the start-to-goal path. We visualize this construction in Fig. 1.
@@ 4.3 The Clever Hans Cheat
To illustrate this in our path-star example, without loss of generality, consider a ground truth path that is of the form $\bm{r}={v}_{\tt{start}},v_{1},v_{2},\ldots,{v}_{\tt{goal}}$. With a slight abuse of the indexing notation, let $\bm{r}_{<i}={v}_{\tt{start}},v_{1},\ldots,v_{i-1}$ be the prefix of length $i$ (so we index from 0 instead of 111). Observe that nodes from $v_{2}$ onwards, until before ${v}_{\tt{goal}}$, have precisely one edge going “away” from ${v}_{\tt{start}}$. Thus, consider when the model is given as input, $(\bm{p},\bm{r}_{<i})$ where $\bm{p}=(\texttt{adj}(G),{v}_{\tt{start}},{v}_{\tt{goal}})$, to fit the target $v_{i}$. The model first merely needs to scan the adjacency list $\texttt{adj}(G)$ within $\bm{p}$ for the one edge containing $v_{i-1}$ in the first position. Then, the model only has to predict the other node on that edge as $v_{i}$. Note though, this cheat cannot work on fitting the target $v_{1}$ given the input $\bm{r}_{<1}={v}_{\tt{start}}$ since ${v}_{\tt{start}}$ has many outward edges — we will scrutinize this node in the next section. We illustrate this difference between $v_{1}$ and the remaining tokens as the “easy” vs. “difficult” tokens in Fig. 2.
Crucially, the above cheating mechanism for fitting the easy tokens does not require any lookahead. It is simple, and implementable by an induction head-like module (Olsson et al., 2022). Then, given the well-known simplicity bias of neural networks (Shah et al., 2020), we hypothesize that the later tokens will be quickly fit and ignored during training. This destroys any gradient signal (a.k.a gradient starvation (Pezeshki et al., 2021)) to efficiently learn the “right-to-left” solution — the solution that requires looking at all the tokens in $\bm{r}$, and learning that they are simply the unique path from ${v}_{\tt{goal}}$ spelled in reverse.
We emphasize two key aspects of this cheating behavior. First, these shortcuts are unlike well-known shortcuts (see Remark 5) that map from the original input prefixes $\bm{p}$ to the ground truth $\bm{r}$. What we identify is unique to the mapping from the teacher-forced prefix $(\bm{p},\bm{r}_{<i})$ to $r_{i}$. Hence, we name this Clever Hans cheating. Another notable point is that this does not come from a dearth of samples: even if we had infinite training data at our disposal, the model can still fit the easy tokens of all that data by Clever Hans cheating.
@@ 5 Experimental Verification
For each experiment, we generate the training and test graphs from the same distribution $\mathcal{D}$, all with the same topology of ${G}_{d,l}(N)$ with fixed $d,l$ and $N$. Thus, any failure we demonstrate is an in-distribution failure, and does not arise from the inability to generalize to different problem lengths (Anil et al., 2022). While the graphs have identical topology, this is not a trivial memorization problem for the model, since the graphs are labeled differently, and the adjacency list randomized — the model has to learn a general algorithm. Throughout the experiments, we fix the number of samples to $200k$ and fix the number of node values to $N=100$ across topologies to enable diverse instantiations of the topology for training and testing.
@@
We evaluate models from two architectural families to highlight that the failures are not tied to a particular architecture but stem from the next-token prediction objective. For Transformers, we use from-scratch GPT-Mini, and pretrained GPT-2 large (Radford et al., 2019). For recurrent models, we use from-scratch Mamba (Gu & Dao, 2023). We optimize using AdamW (Loshchilov & Hutter, 2019) until perfect training accuracy. To rule out grokking behaviour (Power et al., 2022), we trained the cheaper models for as long as 500500500 epochs. More details are in §G.2.
@@
[…]
@@
[…]
Removing the Clever Hans cheat via teacherless training (Tschannen et al., 2023; Monea et al., 2023)
We now consider a training setup where we prevent Clever Hans cheating (Failure 2a) and examine how learning differs. Concretely, consider modifying teacher-forcing by replacing the input $\bm{r}$ (which reveals the ground truth) with an uninformative input $\bm{r}^{\$}$, consisting of the same special (“lookahead”) token \$\\$\$ repeated $l$ times. For supervision in the loss, we still use the original target $\bm{r}$. Thus, the model cannot fit the targets by looking at the prefixes $\bm{r}_{<i}$ and by predicting the next token $v_{i}$ via cheating. Instead, the model only has access to the graph description in $\bm{p}$ to lookahead and fit all the targets $v_{i}$ for $i=1,\ldots,l$. Formally, we maximize:
@@
We denote a model trained this way by $\texttt{LM}_{\theta}^{\$}$ and perform inference simply by conditioning on \$\\$\$ tokens i.e., to extract $\hat{r}_{i}$, we feed the uninformative prefix $\bm{r}^{\$}_{<i}$ as input, rather than autoregressively feeding the output $\hat{\bm{r}}_{<i}$ as input. We denote the resulting accuracy by $\texttt{Acc}_{\texttt{\$}}(\texttt{LM}_{\theta}^{\$})$ (see Eq 9). Our goal however is to evaluate whether forcing the model to lookahead can dodge the Clever Hans cheat, thereby allowing the correct mechanism to be learned.
@@ 6 Related Work
Going beyond next-token prediction. Various works have explored architectures and objectives that train the backbone to go beyond next-token prediction. This includes non-autoregressive models (Gu et al., 2018), energy-based models (Dawid & LeCun, 2023), diffusion models (Gong et al., 2023), and variants of Transformers learning to either predict future tokens in a different ordering (Li et al., 2021) or all at one go (Qi et al., 2020; Monea et al., 2023), or injecting “lookahead” data (Du et al., 2023a).
Teacherless training was proposed for image-to-text captioning models in Tschannen et al. (2023) under the more generic name of parallel prediction with the similar motivation that the tokens in a caption must purely rely on the image rather than parts of the caption itself. Monea et al. (2023) proposed the same idea as “parallel speculative sampling” for the orthogonal goal improving the inference-time compute. On that note, we clarify that research in parallel decoding too is concerned with predicting multiple future tokens (Stern et al., 2018), the goal is purely inference-time efficiency. Finally, it is worth noting that action chunking (Zhao et al., 2023) in imitation learning is a close counterpart of teacherless training although motivated through different arguments.
@@ Acknowledgments
We would like to thank Colin Raffel, Tiago Pimentel, and Surbhi Goel for their extensive feedback on a draft of this preprint, especially for pointers to some key references. We thank Garret Tanzer for pointing us to Tschannen et al. (2023) for a prior instance of the idea of teacherless training.
@@ Appendix B Teacher-Forcing Failure and Snowballing Failure are Distinct
[…]
@@ Appendix D Other Remarks
      Node-ID-agnostic memorization: If the number of training data is $\Omega(|E|!)$, the model can still implement a form of memorization, but this requires a cleverer strategy that is agnostic to the node ID’s. Concretely, for a fixed assignment of node ID’s, there are only $O(|E|!)$ ways to permute the adjacency list. Assume that the model sees all such permutations during training (albeit with varying instantiations of the node IDs). During inference, the model can first look up whether the test input corresponds to an existing permutation it had witnessed (possibly with different node ID’s). For example 12;23;411$\to$ 2;2$\to$ 3;4$\to$ 11 → 2 ; 2 → 3 ; 4 → 1 would correspond to 1020;2030;401010$\to$ 20;20$\to$ 30;40$\to$ 1010 → 20 ; 20 → 30 ; 40 → 10. Then the model merely needs to recall from its memory, the index where the target token is located in the input for that permutation. The model can then look at the same index in the test input and output the node ID located there. Thus, if the target token was 444 for the first sequence above, the model can output 404040 for the other sequence.
@@ E.4 Inference in Teacherless Training
In teacherless training, we make use of an uninformative input $\bm{r}^{\$}$ that simply corresponds to a series of dummy tokens denotes by \$\\$\$. During inference, instead of autoregressing on the model’s own output, we use this uninformative input. We formalize this below:
@@ F.1 Snowball Failure
$\hskip-14.22636pt\text{where }\hskip 5.69054pt\bm{p},\bm{r}\sim\mathcal{D},\hskip 5.69054pt\hat{\bm{r}}_{1<}\stackrel{{\scriptstyle\mathclap{\mbox{{\tiny\tt ag}}}}}{{\sim}}\texttt{LM}_{\theta}(\cdot;\bm{p},r_{1})$
If $\texttt{Acc}_{\texttt{sb}}(\texttt{LM}_{\theta})$ is 1$\approx$ 1≈ 1, then Failure 1 is not prominent in our task. If $\texttt{Acc}_{\texttt{sb}}(\texttt{LM}_{\theta})\ll 1$, then clearly teacher-forcing is responsible for surpressing errors in generation, strongly hinting at the fact that Failure 1 is at play. We display the results in Fig. 5 (left). We observe that the accuracy $\texttt{Acc}_{\texttt{sb}}$ is barely affected even for graphs with very long paths $L=40$.